(a)   A study of 35 gamers showed that their average score on a particular game was 90 and the population standard deviation is 5.

(i)                    Find the best point estimate of the population mean.


 

(ii)                 Find the 95% confidence interval of the mean score for all gamers.




(iii)               Find the 95% confidence interval of the mean score if a sample of 70 gamers is used instead of a sample of 35.


 

(iv)               From your answer in part (ii) and (iii), which interval is smaller?



1
Expert's answer
2021-08-09T11:12:27-0400

(i) The best point estimate for the population mean is the sample mean: "x=90."


(ii) The critical value for "\alpha=0.05"  is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=(x-z_c\times\dfrac{\sigma}{\sqrt{n}}, x+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(90-1.96\times\dfrac{5}{\sqrt{35}}, 90+1.96\times\dfrac{5}{\sqrt{35}})"

"=(88.344, 91.656)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "88.344<\mu<91.656," which indicates that we are 95% confident that the true population mean "\mu"

is contained by the interval "(88.344, 91.656)."


(iii) The critical value for "\alpha=0.05"  is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=(x-z_c\times\dfrac{\sigma}{\sqrt{n}}, x+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(90-1.96\times\dfrac{5}{\sqrt{70}}, 90+1.96\times\dfrac{5}{\sqrt{70}})"

"=(88.829, 91.171)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "88.829<\mu<91.171," which indicates that we are 95% confident that the true population mean "\mu"

is contained by the interval "(88.829, 91.171)."


(iv) The 95% confidence interval for "n=70" is narrower than the 95% confidence interval for "n=35."